OpenStudy (anonymous):

Let f(x)=e^x-e^4x Find all extreme values (if any) of f 0

myininaya (myininaya):

1st step: Find derivative.

OpenStudy (anonymous):

What would that be?

OpenStudy (anonymous):

I did that and then set it equal to 0, but couldnt find what x would equal

myininaya (myininaya):

derivative of e^x is itself.

myininaya (myininaya):

what did you get after differentiating e^(4x)

OpenStudy (anonymous):

i got the whole derivative to be e^x-4e^x and now what do u get when u set that equal to zero?

myininaya (myininaya):

you are missing something

myininaya (myininaya):

\[e^x-4e^{4x}=e^x(1-4e^{3x})\] \[e^x(1-4e^{3x})=0\] Now set both factors =0

OpenStudy (anonymous):

and then what do you get? this is where im stuck

myininaya (myininaya):

Well is e^x ever 0?

OpenStudy (anonymous):

No so thats canceled. Then the other one i got x=-ln(1/4) / 3

OpenStudy (anonymous):

but the answer says it should have it in the form of E

myininaya (myininaya):

(also i don't like to say that cancels; i just there is no solution to that) \[e^{3x}=\frac{1}{4} => \ln(\frac{1}{4})=3x\] did i miss something? how did you get -?

myininaya (myininaya):

You could do ln(1)-ln(4)=-ln(4) I can see how you would the negative then -ln(4)=3x

myininaya (myininaya):

then divide both sides by 3 x=-3ln(4) or x=ln(4^(-3)) or x=ln(1/64) or x=-ln(64)

OpenStudy (anonymous):

I see what you did, now im stuck at ln(1/4)=3x

myininaya (myininaya):

anyways this number doesn't occur on the interval [0,1]

OpenStudy (anonymous):

you divide by 3 and then dont you get ln(1/4)/3

myininaya (myininaya):

or -ln(64)

OpenStudy (anonymous):

why

myininaya (myininaya):

above i put it step by step using properties of log

myininaya (myininaya):

ln(a/b)=ln(a)-ln(b)

myininaya (myininaya):

cln(a)=ln(a^c)

myininaya (myininaya):

ln(ab)=ln(a)+ln(b) ln(1)=0

myininaya (myininaya):

ln(1/4)=ln(1)-ln(4)=0-ln(4)=-ln(4) 1/3*ln(1/4)=1/3* (-ln(4))=-1/3*ln(4)=ln(4^(-1/3)) And I made an error above.

myininaya (myininaya):

but still by properties of log we can rewrite this

OpenStudy (anonymous):

thank you!